OpenStudy (anonymous):

is it possible to make this an equivalent fraction? 3/4=x/3

OpenStudy (anonymous):

\[\frac{3}{4} = \frac{x}{3}\]\[\frac{9}{12} = \frac{4x}{12}\]\[9=4x\]\[\frac{9}{4} = x\] Or you could have just multiplied both sides of the equation by 3:\[(3)\frac{3}{4} = (3)\frac{x}{3}\]\[\frac{9}{4} = x\]

OpenStudy (anonymous):

Thank you!